### v1

# DISCO: Diversifying Sample Condensation for Efficient Model Evaluation

Abstract
Evaluating modern machine learning models has become prohibitively expensive. Benchmarks such as LMMs-Eval and HELM demand thousands of GPU hours per model. Costly evaluation reduces inclusivity, slows the cycle of innovation, and worsens environmental impact. The typical approach follows two steps. First, select an anchor subset of data. Second, train a mapping from the accuracy on this subset to the final test result. The drawback is that anchor selection depends on clustering, which can be complex and sensitive to design choices. We argue that promoting diversity among samples is not essential; what matters is to select samples that maximise diversity in model responses. Our method, Diversifying Sample Condensation (DISCO), selects the top-k samples with the greatest model disagreements. This uses greedy, sample-wise statistics rather than global clustering. The approach is conceptually simpler. From a theoretical view, inter-model disagreement provides an information-theoretically optimal rule for such greedy selection. DISCO shows empirical gains over prior methods, achieving state-of-the-art results in performance prediction across MMLU, Hellaswag, Winogrande, and ARC.
1 Introduction
Figure 1: Imbalance. More evaluation budget is spent on less informative samples in test sets.
[…]
Several efficient evaluation approaches have emerged. A common framework works in two parts: subset selection and performance prediction. The first part selects a static subset of anchor points from the evaluation dataset. The second part predicts full benchmark performance by extrapolating from accuracy on this subset. To select anchor points, existing methods often rely on clustering. Samples are grouped by the similarity of responses they induce in a set of reference models (Vivek et al., 2023; Polo et al., 2024). Variants of this framework include dynamic anchor selection (Hofmann et al., 2025), modified prediction models (Kipnis et al., 2024), and new benchmarks for method comparison (Zhang et al., 2025).
We seek to improve both parts of this framework. For subset selection, we argue that diversity among samples is not essential. What matters is diversity in model responses. We prove that inter-model disagreement is the most informative signal for estimating benchmark performance when the goal is to differentiate and rank models (Proposition 1). Evaluation should therefore focus on samples that elicit varied responses (Figure 1). For performance prediction, we argue that existing methods add unnecessary complexity by estimating hidden model parameters before predicting test performance (Polo et al., 2024; Kipnis et al., 2024). We instead propose a direct route. Model signatures, defined as the concatenation of outputs on the selected subset, serve as inputs to simple predictors of benchmark performance. This framework is simpler, yet matches and surpasses more complex alternatives.
We validate these ideas through Diversifying Sample Condensation (DISCO). DISCO selects a small, informative subset of evaluation samples by focusing on model disagreement. Disagreement is measured by predictive diversity scoring (PDS, Rubinstein et al. (2024)), originally proposed for out-of-distribution detection. A simple metamodel then predicts benchmark performance directly from the model signatures on this subset. We evaluate DISCO in both language and vision domains. On MMLU, for example, DISCO reduces evaluation cost by 99.3% with only 1.07 percentage points of error. Compared with prior methods such as Anchor Points (Vivek et al., 2023), TinyBenchmarks (Polo et al., 2024), and Metabench (Kipnis et al., 2024), DISCO achieves a stronger efficiency–precision trade-off.
2 Related work
We review prior work relevant to our approach. We first highlight the escalating cost of evaluation for contemporary large models and motivate the need for efficiency. We then survey prior attempts at efficient benchmarking, covering instance and task reduction techniques. Finally, we describe our novelty and contributions.
[…]
Label-efficient evaluation. In the pre-LLM context, labelling a test set used to be a cost bottleneck for evaluation. In this context, the concept of “active testing” has been explored, where labelling budget is maximally assigned to information-rich samples (Majumdar & Niksic, 2017; Ji et al., 2021; Deng & Zheng, 2021; Kossen et al., 2021; Hu et al., 2023; Kossen et al., 2022; Huang et al., 2024). In our case, we are concerned with the inference costs of evaluation. As such, active testing approaches are not directly applicable, as they require a full inference over the test set to identify informative samples to label.
Efficient benchmarking. In the LLM era, benchmarks have diversified to measure multiple capabilities and styles of model behaviours. Researchers have proposed strategies to build an efficient benchmark in the first place (Prabhu et al., 2024; Perlitz et al., 2023; Rädsch et al., 2025). There were attempts to compress multiple benchmarks, measuring an array of capabilities of LLMs, into a single one by eliminating redundancies (Zhao et al., 2024; Yuan et al., 2025). Others have focused on selection of small, informative subsets, also known as “Anchor point” approaches (Vivek et al., 2023; Polo et al., 2024; Li et al., 2025; Gupta et al., 2025; Kipnis et al., 2024). Given an entire dataset, they compute a small subset of data points according to the representativeness criterion, determined through the correctness patterns of a large number of source models. Afterwards, the target model performance is estimated based on weighted accuracy computation on the selected subset. In particular, tinyBenchmarks (Polo et al., 2024) and Metabench (Kipnis et al., 2024) have adopted Item Response Theory (IRT) (Lord & Novick, 2008) to estimate model performance in a principled manner. Hofmann et al. (2025) proposed an IRT-based approach to LLM evaluation that selects anchor points dynamically for each model, guided by its predictions on previously chosen anchors. To address the growing number of methods for efficient LLM evaluation, Zhang et al. (2025) recently introduced a large-scale benchmark. In this work, we suggest reconsidering the approaches for both anchor point selection and target model performance estimation, as explained below.
Our novelty and contribution. We differentiate our approach, Diversifying Sample Condensation (DISCO), from previous work in two aspects. (1) model disagreement (Rubinstein et al., 2024) is a simpler and more effective proxy for sample informativeness than representativeness (Vivek et al., 2023; Polo et al., 2024). (2) The application of metamodels on model signatures is a simpler and more effective approach than direct accuracy evaluation approaches (Vivek et al., 2023) or prior approaches that require estimating latent model parameters (Polo et al., 2024; Kipnis et al., 2024).
Figure 2: Problem overview. We aim at selecting a much smaller evaluation dataset than the original evaluation dataset, while keeping the estimated performances as close as possible. Figure 3 details the selection algorithm and the performance predictor.
3 Problem
Our task is the estimation of model performance on a benchmark. Let $f:\mathcal{X}\rightarrow\mathcal{Y}$ be a predictive model over a dataset $\mathcal{D}:=\{(x_{1},y_{1}),\ldots,(x_{N},y_{N})\}$ sampled iid from some distribution. We are interested in estimating the model performance on the dataset $r(f,\mathcal{D})$. An example metric for model performance is accuracy: for a probabilistic classifier $f:\mathcal{X}\rightarrow[0,1]^{C}$, accuracy is defined as $\frac{1}{N}\sum_{i}\mathbf{1}\left(\arg\max_{c}f_{c}(x_{i})=y_{i}\right)$.
We are interested in estimating ${r}(f,\mathcal{D})$ in a cost-effective way. We seek ways to sample a subset of size $K\ll N$ from the original set $\mathcal{D}$ to estimate ${r}(f,\mathcal{D})$. The overall problem is described in Figure 2. An integral ingredient for both prior works and ours is the set of source models $\mathcal{F}=\{f^{1},\ldots,f^{M}\}$, a held-out set of models whose ground-truth performances are known. We define the target models as the models whose performances we aim to estimate.
Figure 3: DISCO overview. First, we select a subset of an evaluation dataset with the most informative samples. Second, we predict the performance of unseen models from their outputs on the selected samples.
4 Solution
This section presents DISCO, our solution to the problem of efficient performance evaluation. DISCO is composed of two steps: (i) the dataset selection, where given an original dataset and an held-out set of source models, we identify a much smaller subset of samples; and (ii) the performance prediction, where given the model outputs on our DISCO selected evaluation set, we estimate the model performance on the original set.
4.1 Dataset selection
At this stage, we require a score that quantifies each sample’s informativeness for predicting performance on the full dataset. Using this score, we rank the samples and select a top-k subset that best preserves the dataset’s information content.
4.1.1 Prior selection methods
We first review existing approaches for selecting representative data points in the evaluation set, referred to as anchor points.
Anchor-conf Vivek et al. (2023) choose $K$ anchors ${A}=\{a_{k}\}_{1}^{K}\subset\{x_{i}\}_{1}^{N}$ that minimise the sum of distances between each data point and the closest anchor: $\min_{A}\sum_{i,k}d\left(e(x_{i}),e(a_{k})\right),$ where the $e(x)$ abbreviates the concatenated model likelihoods $e(x,y):=\left[f^{1}(x)_{y},\ldots,f^{M}(x)_{y}\right]$ for the input and ground-truth label $(x,y)$ for the source models $\mathcal{F}=\{f^{1},\ldots,f^{M}\}$.
Anchor-corr (Polo et al., 2024) is nearly identical to Anchor-conf, except that the embedding uses correctness scores instead of likelihoods: $e(x,y):=\{s^{1}(x,y),\ldots,s^{M}(x,y)\}$, where $s^{m}(x,y):=1(\arg\max_{c}f^{m}(x)_{c}=y)$ encodes correctness of model $f^{m}$ on sample $x$.
Anchor-IRT (Polo et al., 2024) uses the Item-Response Theory (IRT) to define a parametric model $\Pr\left(s^{m}_{i}=1\mid\theta^{m},\alpha_{i},\beta_{i}\right)=\operatorname{sigmoid}(-\alpha_{i}^{\top}\theta^{m}+\beta_{i})$. It predicts the correctness of a model $f^{m}$ on a sample $x_{i}$ with parameters $\theta^{m}\in\mathbb{R}^{d}$, $\alpha_{i}\in\mathbb{R}^{d}$, and $\beta_{i}\in\mathbb{R}$. Using observations of the sample-wise correctness of source models $(x_{i},y_{i},s_{i}^{m})$, the parameters are inferred with an Expectation-Maximisation algorithm. Now, they continue the anchor selection based on the sample-wise embeddings $e(x_{i}):=(\alpha_{i},\beta_{i})$.
Best for validation Kipnis et al. (2024) finds an anchor set $A$ through an iterative search. The algorithm first generates a large number of candidate anchor sets, $\{A_{1},\dots,A_{P}\}$, by uniformly sampling from the full dataset $\mathcal{D}$. For each candidate set $A_{p}$, a simple scalar-to-scalar regression model, $g_{p}$, is trained on the source models $\mathcal{F}$. This model learns to map the performance on the subset, $r(f,A_{p})$, to the known ground-truth performance on the full dataset, $r(f,\mathcal{D})$. Each trained regressor $g_{p}$ is subsequently evaluated on a held-out validation set of models. The final anchor set $A$ is selected as the candidate $A_{p}$ whose corresponding regressor $g_{p}$ yields the lowest prediction error (e.g., RMSE) on this validation set.
How DISCO differs. Unlike clustering, we use sample-wise statistic to determine samples with maximal information content. This greatly simplifies the sampling procedure. We exploit the model diversity, not model confidence or correctness. A set of models can be highly confident and diverse at the same time. We argue that inputs that induce model diversity are more useful for performance prediction.
4.1.2 DISCO selection
We now present our selection method. We hypothesise that samples that produce the greatest diversity in model outputs carry most information about the model performance (see Figure 1). In this part, we explain how we identify such samples in the test dataset. Our sample selection strategies are illustrated in Figure 3. The main approach in Diversifying Sample Condensation (DISCO) is to select a subset $\mathcal{D}_{\text{{DISCO}}}$ of the original evaluation set $\mathcal{D}$ by sampling the top-k samples based on disagreement score, such as Predictive Diversity Score (PDS) from Equation 1. This follows the intuition shown in Figure 1.
We start with an information-theoretic observation below.
Proposition 1.
Let $\mathcal{D}=\{(x_{i},y_{i})\}_{i}^{N}$ be a test set and $m\sim\mathrm{Unif}\{1,\dots,M\}$ be the index of a uniformly chosen model. Let $f_{c}^{m}(x_{i})\in[0,1]$ be the predictive probability for class $c$ of model $f^{m}$ on input $x_{i}$. We write $\widehat{y}_{i}^{m}$ for the categorical random variable following $\operatorname{Cat}(f_{1}^{m}(x_{i}),\ldots,f_{C}^{m}(x_{i}))$. Define ensemble mean prediction to be $\widebar{f}_{c}(x_{i}):=\mathbb{E}_{m}[f^{m}_{c}(x_{i})]$ for each class $c$ and define corresponding prediction random variable as $\widehat{y}_{i}$ following $\operatorname{Cat}(\widebar{f}_{1}(x_{i}),\ldots,\widebar{f}_{C}(x_{i}))$. Let $S(m)=S(f^{m},\mathcal{D})$ denote a function of model $m$ and dataset $\mathcal{D}$, such as model accuracy, that is injective with respect to $m$. We make additional Assumptions A1A2 in appendix. Then,
$\mathrm{MI}_{m,\widehat{y}_{i}}\left(S(m);\widehat{y}_{i}\right)\;=\;{H}(\widehat{y}_{i})-\mathbb{E}_{m}\left[{H}\left(\widehat{y}^{m}_{i}\right)\right]\;=\;\mathrm{JSD}\left(\widehat{y}_{i}^{1},\ldots,\widehat{y}_{i}^{M}\right).$
where ${H}(\cdot)$ is entropy, $\mathrm{MI}(\cdot)$ is mutual information, and $\operatorname{JSD}(\cdot)$ is generalised Jensen-Shannon Divergence for multiple distributions (Fuglede & Topsoe, 2004).
See proof in Appendix A.
We conclude that the sample $i$ conveying the greatest level of information for the prediction of $S(m)$ (e.g. model accuracy) is the one with greatest $\operatorname{JSD}\left(\widehat{y}_{i}^{1},\ldots,\widehat{y}_{i}^{M}\right)$. This generalised Jensen-Shannon divergence translates to the diversity of distributions (Fuglede & Topsoe, 2004). Based on the insight that model diversity matters for performance prediction, we also consider an alternative measure that measures the model diversity: predictive diversity score (PDS) (Rubinstein et al., 2024). It is more interpretable, as it is a continuous generalisation of the number of unique argmax category predictions among $M$ source models:
$\mathrm{PDS}\left(\widehat{y}_{i}^{1},\ldots,\widehat{y}_{i}^{M}\right):=\frac{1}{C}\sum_{c}\max_{m}f_{c}^{m}(x_{i}).$ (1)
PDS is related to JSD through the enveloping inequalities below:
Proposition 2.
Denoting $\mathrm{PDS}_{i}:=\mathrm{PDS}\left(\widehat{y}_{i}^{1},\ldots,\widehat{y}_{i}^{M}\right),\mathrm{JSD}_{i}:=\mathrm{JSD}\left(\widehat{y}_{i}^{1},\ldots,\widehat{y}_{i}^{M}\right)$ for each sample $i$, we have
$\frac{2}{M^{2}\ln 2}\,(\mathrm{PDS}_{i}-1)^{2}\;\leq\;\mathrm{JSD}_{i}\!\;\leq\;\frac{M}{M-1}\log M\cdot(\mathrm{PDS}_{i}-1).$
See proof in Appendix B.3.
In the experiments, we consider both JSD and PDS as criteria for sample selection.
4.2 Performance prediction
Once a subset of dataset samples $A$ is selected, we use the responses of the target model $f$ on $A$ to estimate the true performance.
4.2.1 Prior prediction methods
We first review existing approaches for estimating the true performance using predictions on anchor points ${A}=\{a_{1},\ldots,a_{K}\}$.
Weighted sum Vivek et al. (2023) estimates the true performance by directly computing the accuracy on the anchor set: $\operatorname{WS}(f,A):=({1}/K)\sum_{k}w_{k}\,s^{m}_{k}$, where $w_{k}$ is the number of original training samples $x_{i}$ assigned to the anchor $a_{k}$ in the Anchor-Corr method.
p-IRT (Polo et al., 2024): makes adjustments to the vanilla accuracy on the anchor set by adding a correction term derived from the IRT in Anchor-IRT in: $\operatorname{p-IRT}(f,A):=(1/K)\sum_{k\in A}s_{k}+1/(N-K)\sum_{k\notin A}p_{i}$, where $\hat{p}_{i}$ is the IRT estimation computed based on the parameters obtained in Anchor-IRT.
gp-IRT (Polo et al., 2024) is a mixture of the two approaches above: $\operatorname{gp-IRT}(f,A)=\lambda\cdot\operatorname{WS}(f,A)+(1-\lambda)\cdot\operatorname{p-IRT}(f,A)$ where $\lambda\in[0,1]$.
ability-IRT Kipnis et al. (2024) is a two-stage method that uses the anchor set $A$ as a diagnostic tool rather than just a miniature test. First, it uses a pre-calibrated IRT model to estimate a latent “ability” score, $\hat{\theta}^{f}$, from the target model’s pattern of correct and incorrect responses on $A$. Second, a pre-trained regressor, $g$, predicts the final performance $S^{f}$ using both the simple anchor set accuracy $\hat{S}^{f}_{A}$ and this more informative ability score $\hat{\theta}^{f}$ as input features. The final prediction is given by $S^{f}=g(\hat{S}^{f}_{A},\hat{\theta}^{f})$, leveraging a deeper measure of the model’s capability to improve the estimate.
How DISCO differs. Previous prediction methods rely on scalar summaries of performance, such as the (weighted or corrected) accuracy on the anchor set. In contrast, our approach leverages a much richer signal: the model signature, defined as the concatenation of the model’s raw outputs on the selected samples. By learning a direct mapping from the high-dimensional signature to the final performance, we bypass the complexities of psychometric modeling and demonstrate that a simpler, more direct approach can be more effective.
4.2.2 DISCO prediction
Given a smaller set of test dataset $\mathcal{D}_{\text{{DISCO}}}$, we estimate the performance of a model $f$ as closely as possible to the true full test performance ${r}(f,\mathcal{D})$. We deliberately opt for simple approaches here, in order to make a point that simple is best; we also compare against a rather complex prior work and show that our simple method wins against the stolen. Our performance prediction framework is depicted in Figure 3.
Model signatures. We hypothesise that models with similar output patterns on $\mathcal{D}_{\text{{DISCO}}}$ will exhibit similar performance. To capture this pattern, we define a model signature as the concatenation of the model’s outputs on $\mathcal{D}_{\text{{DISCO}}}$: $f(\mathcal{D}_{\text{{DISCO}}}):=\left[f(x_{1}),\ldots,f(x_{L})\right]$.
Such function signature may have a large dimensionality, as it is the product of model output dimensionality (e.g. 1000 for ImageNet) and the number of selected samples $|\mathcal{D_{\text{DISCO}}}|$ (e.g. can go up to 50k for ImageNet validation set). To reduce the storage burden and improve generalizability, we consider applying a dimenionality reduction technique based on principal component analysis (PCA): $Q\circ{f}(\mathcal{D}_{\text{{DISCO}}})$.
KNN prediction. Built on the hypothesis that the similarities in function signature imply performance similarity, we consider the kNN predictor based on a held-out set of models $\mathcal{F}$. Given a function $f$ to evaluate, we identify the K most similar models in $\mathcal{F}$ using the Euclidean distance between their signatures after dimensionality reduction. We estimate $f$’s performance by averaging the performances of the K most similar models.
Parametric mapping. We also consider a parametric prediction variant. A single parametric mapping $R$ is trained for the prediction of model performance. As the training set, we use $M$ model signatures $Q\circ{f}_{1}(\mathcal{D}_{\text{{DISCO}}}),\ldots Q\circ{f}_{M}(\mathcal{D}_{\text{{DISCO}}})$ for $\mathcal{F}$ as the training set for the regression problem of training a mapping $R(\cdot)$ to let $R\circ Q\circ{f}_{m}(\mathcal{D}_{\text{{DISCO}}})$ approximate $\widehat{r}(f,\mathcal{D})$. The predictor $R$ can be implemented using a neural network, linear regression, or a Random Forest, for example.
5 Experiments
In this section, we introduce the experimental setup, present the main results of Diversifying Sample Condensation (DISCO) in language domain (§5.1), analyse contributing factors (§5.2), and demonstrate that the method is domain-agnostic and can also be successfully applied to the vision domain (§5.3).
Datasets. We evaluate DISCO on two widely used language modeling benchmarks: MMLU (Hendrycks et al., 2021) and HellaSwag (Zellers et al., 2019). MMLU covers 57 tasks about world knowledge and problem-solving ability, and HellaSwag focuses on commonsense natural language inference.
Models. Building on the TinyBenchmarks framework (Polo et al., 2024), we evaluate 424 large language models (LLMs) from Hugging Face’s Open LLM Leaderboard (Fourrier et al., 2024). The models cover GPT- (Radford et al., 2019), LLaMA- (Touvron et al., 2023), DeepSeek- (DeepSeek-AI et al., 2025), and BERT-style (Devlin et al., 2019) architectures, with model sizes ranging from 1.3 billion to 32 billion parameters.
Model split. DISCO is based on a meta-model approach where a predictor is constructed based on the model signatures of a pool of source models $\mathcal{F}$ and tested on a disjoint set of models. This approach has traditionally been criticised for its dependency on the set of existing models: the approach may fail to retain performance with unforeseen changes in future models. To address this concern, we introduce the chronological split, where the source models $\mathcal{F}$ consist of models published before January 13, 2024 and the meta test set consists of models after the cutoff date. The train-test ratio is 9:1.
Evaluation. We evaluate DISCO and baseline approaches using two complementary metrics. First, the Mean Absolute Error (MAE) of the model accuracies, reported as percentage points (%p), captures the absolute error of accuracy prediction. Second, to assess the consistency of the relative ordering of models, we report the Spearman rank correlation (Rank) in model ranking between the true and estimated model performances.
5.1 Main results
Approach Selection Prediction MMLU (14k) HS (10k) WG (1.3k) ARC (1.2k)
[…]
Baseline Random Direct eval. 3.45 0.916 2.85 0.839 3.60 0.827 2.61 0.898
tinyBenchmarks Random gp-IRT 2.79 0.922 1.96 0.819 1.64 0.928 2.22 0.921
Anchor-IRT gp-IRT 3.25 0.922 2.19 0.830 2.24 0.850 4.55 0.708
Anchor-corr gp-IRT 2.08 0.927 1.27 0.937 1.95 0.918 2.18 0.948
Metabench Best for val. ability-IRT 2.08† 0.904† 0.80 0.974† 1.23† 0.947† 1.14 0.971
Model signature Random Sig. + kNN 1.82 0.912 1.49 0.899 1.58 0.920 2.30 0.905
Sig. + RF 1.81 0.933 1.36 0.938 1.29 0.926 1.72 0.938
DISCO (ours) High PDS Sig. + kNN 1.31 0.972 1.32 0.956 1.19 0.951 1.96 0.937
Sig. + RF 1.07 0.987 1.01 0.984 1.00 0.967 1.47 0.971
High JSD Sig. + kNN 1.14 0.975 1.50 0.944 1.26 0.955 2.11 0.939
Sig. + RF 1.30 0.987 0.86 0.972 1.09 0.973 1.75 0.938
Table 1: DISCO achieves state-of-the-art test-set compression by using model signatures combined with PDS for accurate performance prediction. Compression of MMLU, HellaSwag (HS), Winogrande (WG) and ARC datasets by DISCO (ours), tinyBenchmarks, Metabench, and other baselines. For each dataset, we reduce the test set to 100 data points (except for Metabench, see below), achieving inference cost reduction of 99.3% and 99.0%, on MMLU and HS, respectively. Sig. + RF/kNN stands for model signature with Random Forest/kNN prediction (§ 4.2.2). Mean absolute error (MAE) is the %p difference in accuracy, and Rank is the Spearman rank correlation between the true model ranking and the estimated model ranking.
†  Results for Metabench are not directly comparable, as it requires more examples to converge: 150 datapoints for MMLU and ARC (+50%), 450 for HS (+350%), and 200 for WG (+100%).
Table 1 shows the main results. Uniform random sampling, together with direct evaluation with the corresponding annotated labels, yields 3.45%p MAE and .916 rank correlation at 100 samples. The approaches introduced in tinyBenchmarks Polo et al. (2024) improve over this baseline, confirming their findings.
We measure the efficacy of DISCO in two steps: adopt model-signature approach on top of uniform random sample selections first, and then consider sampling according to predictive diversity scoring (PDS). Even without PDS, on uniform random samples, model signatures are achieving 1.81%p MAE and .933 rank correlation with Random Forest (RF), reaching the state-of-the-art performance with simple and practical ingredients. When PDS is further considered for sample selection, to diversify the model outputs, we achieve 1.07%p MAE and .987 rank correlation, demonstrating a significant leap from the prior state of the art from tinyBenchmarks Polo et al. (2024) from ICML 2024.
Figure 4: True and estimated performance on MMLU. Scatter plot of performances of 40 models.
To provide an understanding of the distributional comparison of the true model performances and the estimated performances, we show a scatter plot in Figure 4. As signified by the high Pearson’s correlation coefficient at .986, the estimated performances closely follow the true performances.
Figure 5: MMLU performance estimation vs. compression rates. Mean absolute error (MAE), measured in %p difference in accuracy, and the Spearman rank correlation between the true model ranking and the estimated model ranking are shown. At 100 samples, the results are identical to Table 1. Main observations: DISCO hits a better efficiency-precision trade-off across all range of compression rates. For extreme compression rate, kNN is a better choice than random forest (RF).
Figure 5 shows the performance against varying degrees of the test set reduction. We observe that the ranking of estimated evaluation methodologies does not change much across a wide range of degrees of reduction. In particular, our DISCO is consistently the best method across all ranges of number of samples involved. For the extreme rates of compression, at 10 samples, the non-parametric performance predictor of kNN yields better performance than the parametric Random Forest, suggesting that non-parametric approaches may be more suitable at extreme compression.
5.2 Factors analysis
We analyse the impact of several design choices involved in our DISCO on MMLU dataset. See Table 2 for an overview.
Model split. In a recent benchmark for efficient LLM evaluation Zhang et al. (2025), the authors observed that prediction performance drops sharply when test models outperform training models. We extend this idea by replacing performance-based splits with chronological splits, training on older models and testing on newer ones. This better reflects real-world usage, whereas performance-based splits create an artificial stress test.
For this purpose, we introduced the chronological split in §5. We examine the impact of this model splitting on the result. We observe that our DISCO is robust to the choice of splitting strategy. Chronological splitting yields a rank correlation of .987, which is nearly identical to the .986 obtained with uniform splitting (Table 2 (a)).
Table 2: Factor analysis for DISCO on MMLU. Highlighted in bold are the default design choices for DISCO. All comparisons are based on 100 selected samples.
Stratification. We measure the efficacy of the stratification strategy in (Polo et al., 2024) where equal numbers of anchor points are selected from each of 57 tasks in MMLU dataset (Table 2 (b)). We find that stratification (.978) is not effective when data points are sampled according to PDS (.987).
Number of source models. We analyse the sensitivity of DISCO to the number of source models $|\mathcal{F}|$ (Table 2 (c)). With only 100 models (.969 rank correlation), it already outperforms TinyBenchmarks, which uses all 382 available source models (.927 in Table 1). As the number of source models increases, rank correlation steadily improves, reaching a maximum of .987 for $|\mathcal{F}|=382$.
Dimensionality reduction. We compare PCA with different target dimensions to Uniform Manifold Approximation and Projection (UMAP) (McInnes et al., 2020) for dimensionality reduction (Table 2 (d)). We notice that dimensionality reduction helps reduce potential overfitting: without it (using all 3100 dimensions), the correlation is .918, while with PCA at 256 dimensions, it improves to .987. Overall, PCA outperforms UMAP and remains robust across a wide range of dimensions.
Prediction model. We compare consider a wide range of prediction models (Table 2 (e)). Random Forest achieves the highest rank correlation of .987, outperforming all other methods.
5.3 Results for vision domain
Approach Selection Prediction IN val (50k)
[…]
Baseline Random Direct eval. 3.03 0.652
Model signature Random Sig. + kNN 1.72 0.808
Sig. + RF 0.86 0.944
DISCO (ours) High PDS Sig. + kNN 1.68 0.819
Sig. + RF 0.63 0.969
Table 3: DISCO compression of ImageNet validation dataset. We evaluate the generalisation of our DISCO to the computer vision domain. The main metrics are mean absolute error (MAE), measured in %p difference in accuracy, and the Spearman rank correlation (Rank) between the true model ranking and the estimated model ranking. Main observations: (1) Same as for language experiments, model signature is an effective strategy for performance estimation. (2) Using PDS on top improves performance even more.
In this section, we give a quick overview of the DISCO applied to the vision domain. For detailed results, see § C.
Setup. We use ImageNet-1k (Russakovsky et al., 2015) with 1.28M images and 400 pretrained models from timm (Wightman, 2019), spanning convolutional (Krizhevsky et al., 2012) and transformer (Dosovitskiy et al., 2021) architectures (0.3M–300M parameters). Following the language domain, we adopt a chronological split with cutoff 5 April 2023 (88:12 train–test). Performance is evaluated using mean absolute error (MAE) and Spearman rank correlation.
Results. Our DISCO approach significantly compresses the ImageNet validation set by reducing it to just 100 data points, achieving an inference cost reduction of 99.8%. DISCO with uniform random sampling and random forest prediction on model signatures achieve 0.86%p MAE and .944 rank correlation, surpassing the baseline. Using a predictive diversity score (PDS) for data selection and a Random Forest for prediction, our method achieves a 0.63%p MAE and .969 rank correlation, substantially outperforming the baseline (Table 3). The results demonstrate that DISCO is effective in both language and vision domains.
6 Conclusion
Evaluating ML models is increasingly expensive due to larger models, datasets, and benchmarks. It is especially true for general-purpose LLMs requiring broad evaluation.
We propose DISCO, which selects a small informative subset of the evaluation data that maximizes LLM disagreement and estimates model performance from predictions on it. DISCO cuts evaluation costs by over 99% with minimal error and consistently outperforms prior methods.
This enables practical use: efficient evaluation on limited compute, frequent performance tracking during training, and cheap end-user checks of deployed models.
Limitations. The main limitation of DISCO is robustness to distribution shifts in the model population. Shifts can arise from new architectures, training methods, or objectives, introducing patterns unseen during training and reducing estimator accuracy. Future work could address this with adaptive sample selection or periodic retraining on newer models.
Author Contributions
Benjamin, Joon, and Alexander conceived the project. Alexander led the language experiments, Benjamin led the vision experiments. Joon and Martin helped design the experiments. Alexander, Martin, and Joon led the writing of the paper. Martin and Joon provided helpful feedback throughout the project.
Acknowledgments
This work was supported by the Tübingen AI Center. AR thanks the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for support. This research utilized compute resources at the Tübingen Machine Learning Cloud, DFG FKZ INST 37/1057-1 FUGG.
Appendix A Mutual Information and Jensen-Shannon Divergence
In this section, we show that Mutual Information is equivalent to JSD in our setting. We present the setup and assumptions, then prove the proposition.
A.0.1 Setup.
Denote a set of source models as $\mathcal{F}=\{f^{1},\ldots,f^{M}\}$. Let $m\sim\mathrm{Unif}\{1,\dots,M\}$ be the index of a uniformly chosen model and $S(m)=S(f^{m},\mathcal{D})$ denote any statistic that is a deterministic function of $m$ computed on $\mathcal{D}$ (e.g. accuracy on $\mathcal{D}$). Let $f_{c}^{m}(x_{i})\in[0,1]$ be the predictive probability for class $c$ of model $f^{m}$ on input $x_{i}$. We write $\widehat{y}_{i}^{m}$ for the categorical random variable following $\operatorname{Cat}(f_{1}^{m}(x_{i}),\ldots,f_{C}^{m}(x_{i}))$ (i.e., given $m$, the prediction on datapoint $i$ is a categorical random variable $\widehat{y}_{i}^{m}$). Define ensemble mean prediction to be $\widebar{f}_{c}(x_{i}):=\mathbb{E}_{m}[f^{m}_{c}(x_{i})]$ for each class $c$ and define corresponding prediction random variable as $\widehat{y}_{i}$ following $\operatorname{Cat}(\widebar{f}_{1}(x_{i}),\ldots,\widebar{f}_{C}(x_{i}))$.
As a consequence, the ensemble mean prediction random variable $\widehat{y}_{i}$ coincides with marginal distribution of $\widehat{y}_{i}^{m}$:
$\Pr(\widehat{y}_{i}=c)\;=\;\mathbb{E}_{m\sim\operatorname{Unif}[M]}\!\left[\Pr(\widehat{y}_{i}=c|m)\right]\;=\;\frac{1}{M}\sum_{m=1}^{M}\Pr(\widehat{y}_{i}^{m}=c).$
Definition 1 (Jensen–Shannon divergence).
Generalised Jensen-Shannon Divergence for multiple distributions (Fuglede & Topsoe, 2004):
$\mathrm{JSD}\!\big(\widehat{y}_{i}^{1},\ldots,\widehat{y}_{i}^{M}\big)=\frac{1}{M}\sum_{m=1}^{M}D_{\mathrm{KL}}(\widehat{y}_{i}^{m}\|\widehat{y}_{i})=H(\widehat{y}_{i})-\tfrac{1}{M}\sum_{m=1}^{M}H(\widehat{y}_{i}^{m}).$
A.0.2 Assumptions.
Assumption A1 (Deterministic predictions).
Conditional on $m$, each prediction $\widehat{y}_{i}^{m}$ is fully determined by $m$ (or more generally, any residual randomness is independent across $i$ and independent of $m$).
Assumption A2 (Uniform prior).
The model index is uniformly distributed: $m\sim\mathrm{Unif}\{1,\dots,M\}$.
A.0.3 Proposition.
Proposition 3.
Under Assumptions A1A2, if $S(m)$ is injective, then
$\mathrm{MI}_{m,\widehat{y}_{i}}\left(S(m);\widehat{y}_{i}\right)\;=\;{H}(\widehat{y}_{i})\;-\tfrac{1}{M}\sum_{m=1}^{M}H(\widehat{y}_{i}^{m})\;=:\;\mathrm{JSD}\left(\widehat{y}_{i}^{1},\ldots,\widehat{y}_{i}^{M}\right).$
where ${H}(\cdot)$ is entropy, $\mathrm{MI}(\cdot)$ is mutual information.
Proof.
By Assumption A1 and since $S(m)$ is a deterministic function of $m$, we have the Markov chain
$\widehat{y}_{i}^{m}\;\longleftrightarrow\;m\;\longleftrightarrow\;S(m).$
If $S$ is injective, then $m$ is recoverable from $S(m)$, hence
$\mathrm{MI}\big(S(m);\widehat{y}_{i}\big)=\mathrm{MI}\big(m;\widehat{y}_{i}\big).$
By the definition of mutual information,
$\mathrm{MI}(m;\widehat{y}_{i})={H}(\widehat{y}_{i})-\mathbb{E}_{m}\!\left[{H}\left(\widehat{y}_{i}|m\right)\right].$
Conditional entropy (using Assumption A2):
$\mathbb{E}_{m}\big[{H}(\widehat{y}_{i}|m)\big]=\tfrac{1}{M}\sum_{m=1}^{M}{H}(\widehat{y}_{i}^{m}).$
Combine:
$\mathrm{MI}(m;\widehat{y}_{i})={H}(\widehat{y}_{i})-\tfrac{1}{M}\sum_{m=1}^{M}{H}(\widehat{y}_{i}^{m})=:\mathrm{JSD}\left(\widehat{y}_{i}^{1},\ldots,\widehat{y}_{i}^{M}\right).$
Appendix B Bounds for Jensen-Shannon Divergence (JSD) via Predictive Diversity Score (PDS)
In this section, we show that JSD (Definition 1) is bounded quadratically below and linearly above by PDS. We first relate JSD to total variation (§ B.1), then show total variation is monotone in PDS (§ B.2), and then combine these results in § B.3.
B.1 Bounds for JSD via Total Variation (TV)
We begin by showing that JSD is bounded quadratically below and linearly above by total variation. We first introduce the setup with required definitions (§ B.1.1), then prove the proposition (§ B.1.2).
B.1.1 Setup.
Use notation from Appendix A. Let $\widehat{y}_{i}^{m}$ for $m=1,\ldots,M$ be categorical random variables on $K$ classes, and let $\widehat{y}_{i}$ be the ensemble mean prediction random variable defined as the mixture
$\Pr(\widehat{y}_{i}=c)\;=\;\frac{1}{M}\sum_{m=1}^{M}\Pr(\widehat{y}_{i}^{m}=c).$
Definition 2 (Total variation).
For categorical random variables $X,Y$ on the same support,
$\mathrm{TV}(X,Y)=\tfrac{1}{2}\sum_{c}|\Pr(X=c)-\Pr(Y=c)|.$
B.1.2 Proposition.
Now, we show that JSD is bounded quadratically below and linearly above by total variation.
Proposition 4 (JSD–TV sandwich bounds).
For any $M\geq 2$ categorical random variables $\{\widehat{y}_{i}^{m}\}_{m=1}^{M}$ with ensemble mean $\widehat{y}_{i}$,
$\frac{2}{\ln 2}\cdot\frac{1}{M}\sum_{m=1}^{M}\mathrm{TV}(\widehat{y}_{i}^{m},\widehat{y}_{i})^{2}\;\leq\;\mathrm{JSD}\!\big(\widehat{y}_{i}^{1},\ldots,\widehat{y}_{i}^{M}\big)\;\leq\;\frac{M}{M-1}\,\log M\cdot\frac{1}{M}\sum_{m=1}^{M}\mathrm{TV}(\widehat{y}_{i}^{m},\widehat{y}_{i}).$
Proof.
Lower bound. From Definition 1,
$\mathrm{JSD}\!\big(\widehat{y}_{i}^{1},\ldots,\widehat{y}_{i}^{M}\big)=\frac{1}{M}\sum_{m=1}^{M}D_{\mathrm{KL}}(\widehat{y}_{i}^{m}\|\widehat{y}_{i}).$
By Pinsker’s inequality (e.g. Equation 1 in (Sason, 2015)),
$D_{\mathrm{KL}}(X\|Y)\;\geq\;\frac{2}{\ln 2}\,\mathrm{TV}(X,Y)^{2}.$
Substituting $X=\widehat{y}_{i}^{m}$, $Y=\widehat{y}_{i}$ and averaging over $m$ yields the lower bound.
Upper bound. Fix $m$. For each class $c$, write
$\Pr(\widehat{y}_{i}=c)=\alpha\Pr(\widehat{y}_{i}^{m}=c)+(1-\alpha)\zeta(c),\quad\alpha=\tfrac{1}{M},\quad\zeta(c)=\tfrac{1}{M-1}\sum_{s\neq m}\Pr(\widehat{y}_{i}^{s}=c).$
Define $t(c)=\zeta(c)/\Pr(\widehat{y}_{i}^{m}=c)$ when $\Pr(\widehat{y}_{i}^{m}=c)>0$ (set $t(c)=+\infty$ if $\Pr(\widehat{y}_{i}^{m}=c)=0,\zeta(c)>0$). Then $\mathbb{E}_{\widehat{y}_{i}^{m}}[t]=1$ and
$D_{\mathrm{KL}}(\widehat{y}_{i}^{m}\|\widehat{y}_{i})=\mathbb{E}_{\widehat{y}_{i}^{m}}\!\left[-\log\big(\alpha+(1-\alpha)t\big)\right].$
Let $g(u)=-\log(\alpha+(1-\alpha)u)$, $u\geq 0$. Then $g$ is convex, decreasing, with $g(1)=0$, $g(0)=\log(1/\alpha)=\log M$. By convexity,
$g(u)\;\leq\;(1-u)g(0)=(1-u)\log M.$
Thus,
$D_{\mathrm{KL}}(\widehat{y}_{i}^{m}\|\widehat{y}_{i})\leq\log M\cdot\mathbb{E}_{\widehat{y}_{i}^{m}}\!\big[(1-t)\big]\leq\log M\cdot\mathbb{E}_{\widehat{y}_{i}^{m}}\!\big[(1-t)_{+}\big].$
Now
$\mathbb{E}_{\widehat{y}_{i}^{m}}[(1-t)_{+}]=\sum_{c}\Pr(\widehat{y}_{i}^{m}=c)\max\{0,1-\zeta(c)/\Pr(\widehat{y}_{i}^{m}=c)\}=\sum_{c}(\Pr(\widehat{y}_{i}^{m}=c)-\zeta(c))_{+}.$
By the balance-of-deviations identity (§ 1),
$\sum_{c}(\Pr(\widehat{y}_{i}^{m}=c)-\zeta(c))_{+}=\mathrm{TV}(\widehat{y}_{i}^{m},\zeta).$
Finally, since $\Pr(\widehat{y}_{i}=c)=\alpha\Pr(\widehat{y}_{i}^{m}=c)+(1-\alpha)\zeta(c)$, one has
$\mathrm{TV}(\widehat{y}_{i}^{m},\zeta)=\tfrac{M}{M-1}\,\mathrm{TV}(\widehat{y}_{i}^{m},\widehat{y}_{i}).$
Combining yields
$D_{\mathrm{KL}}(\widehat{y}_{i}^{m}\|\widehat{y}_{i})\;\leq\;\tfrac{M}{M-1}\,\log M\cdot\mathrm{TV}(\widehat{y}_{i}^{m},\widehat{y}_{i}).$
Averaging over $m$ gives the upper bound. ∎
Remark 1.
The lower bound is quadratic in total variation, the upper bound linear. Thus, JSD interpolates between quadratic growth near equality and linear growth in worst-case separation.
B.2 Bounds for Total Variation via Predictive Diversity Score
We next show that total variation is monotone in PDS. We introduce the setup with definitions and lemmas (§ B.2.1), then prove the proposition (§ B.2.2).
B.2.1 Setup.
Fix a class $c$. Let $X_{m}=\Pr(\widehat{y}_{i}^{m}=c)$ and $\mu=\Pr(\widehat{y}_{i}=c)$. Define:
Definition 3 (Envelope and spread, per class).
$E_{c}=\max_{m}X_{m}-\mu,\qquad U_{c}=\frac{1}{2M}\sum_{m=1}^{M}|X_{m}-\mu|.$
Definition 4 (Predictive Diversity Score).
$\operatorname{PDS}\!\big(\widehat{y}_{i}^{1},\ldots,\widehat{y}_{i}^{M}\big)=\sum^{K}_{c=1}\max_{m}\Pr(\widehat{y}_{i}^{m}=c).$
Lemma 1 (Balance-of-deviations identity).
For any $a_{1},\dots,a_{M}$ with $\sum_{m}a_{m}=0$, writing $a_{+}=\max\{0,a\}$,
$\sum_{m=1}^{M}(a_{m})_{+}=\sum_{m=1}^{M}(a_{m})_{-}=\tfrac{1}{2}\sum_{m=1}^{M}|a_{m}|.$
Proof.
Decompose $a=a_{+}-a_{-}$, $|a|=a_{+}+a_{-}$. Summing and using $\sum_{m}a_{m}=0$ gives
$\sum_{m}a_{m,+}-\sum_{m}a_{m,-}=0\quad\Rightarrow\quad\sum_{m}a_{m,+}=\sum_{m}a_{m,-}.$
Then
$\sum_{m}|a_{m}|=\sum_{m}(a_{m,+}+a_{m,-})=2\sum_{m}a_{m,+}.$
Applying Lemma 1 with $a_{m}=X_{m}-\mu$ yields
$U_{c}=\tfrac{1}{M}\sum_{m:X_{m}>\mu}(X_{m}-\mu).$
B.2.2 Proposition.
Now, we show that total variation is monotone in PDS.
Proposition 5 (Spread–envelope bounds).
Use notation from Appendix B.2.1. For each class $c$, if at most $z$ models satisfy $X_{m}>\mu$, then
$\frac{1}{M}E_{c}\;\leq\;U_{c}\;\leq\;\frac{z}{M}E_{c}.$
Aggregating over classes,
$\frac{1}{M}E\;\leq\;U\;\leq\;\frac{z}{M}E,$
where
$E=\sum_{c=1}^{K}E_{c},\qquad U=\sum_{c=1}^{K}U_{c}=\tfrac{1}{M}\sum_{m=1}^{M}\mathrm{TV}(\widehat{y}_{i}^{m},\widehat{y}_{i}).$
Proof.
If $E_{c}=0$, then $X_{m}=\mu$ for all $m$ so $U_{c}=0$. Otherwise, let $m^{\star}=\arg\max_{m}X_{m}$. Then
$U_{c}=\tfrac{1}{M}\sum_{m:X_{m}>\mu}(X_{m}-\mu)\;\geq\;\tfrac{1}{M}(X_{m^{\star}}-\mu)=\tfrac{1}{M}E_{c}.$
For the upper bound, each positive term is at most $E_{c}$, and there are at most $z$ such terms, hence
$U_{c}\leq\tfrac{z}{M}E_{c}.$
Summing over classes gives the aggregated bound. ∎
B.3 Final Sandwich Inequality
Finally, we combine results from § B.1.1 and § B.2 to show that JSD is bounded quadratically below and linearly above by PDS.
Proposition 6 (JSD–PDS sandwich).
Use notation from Proposition 1.
$\frac{2}{M^{2}\ln 2}\,(\mathrm{PDS}\big(\widehat{y}_{i}^{1},\ldots,\widehat{y}_{i}^{M}\big)-1)^{2}\;\leq\;\mathrm{JSD}\!\big(\widehat{y}_{i}^{1},\ldots,\widehat{y}_{i}^{M}\big)\;\leq\;\frac{M}{M-1}\log M\cdot(\mathrm{PDS}\big(\widehat{y}_{i}^{1},\ldots,\widehat{y}_{i}^{M}\big)-1).$
Proof.
From Theorem 4,
$\mathrm{JSD}\;\geq\;\tfrac{2}{\ln 2}\,U^{2},\qquad\mathrm{JSD}\;\leq\;\tfrac{M}{M-1}\,\log M\cdot U.$
By the definitions,
$U\;=\;\sum_{c=1}^{K}\frac{1}{2M}\sum_{m=1}^{M}|\Pr(\widehat{y}_{i}^{m}=c)-\Pr(\widehat{y}_{i}=c)|=\frac{1}{M}\sum_{m=1}^{M}\mathrm{TV}(\widehat{y}_{i}^{m},\widehat{y}_{i}),$
and
$E\;=\;\sum_{c=1}^{K}\Big(\max_{m}\Pr(\widehat{y}_{i}^{m}=c)-\Pr(\widehat{y}_{i}=c)\Big)=\mathrm{PDS}\big(\widehat{y}_{i}^{1},\ldots,\widehat{y}_{i}^{M}\big)-1.$
From Proposition 5,
$\tfrac{1}{M}(\mathrm{PDS}-1)\leq U\leq\tfrac{z}{M}(\mathrm{PDS}-1).$
Combining and noticing that $1\leq z\leq M$ yields the quadratic lower bound and linear upper bound in $(\mathrm{PDS}-1)$. ∎
Appendix C Vision results
We introduce the setup and present results.
Dataset. We use ImageNet-1k (Russakovsky et al., 2015) with 1.28 million images. Models. We consider 400 models from timm (Wightman, 2019) that are pretrained on ImageNet. The models cover convolutional (Krizhevsky et al., 2012) and transformer (Dosovitskiy et al., 2021) architectures. Model sizes range from 0.3M to 300M parameters.
Model Split. As in the language domain, we use the chronological split. The cutoff date is 5 April 2023. The train-test ratio of models is 88:12.
Evaluation. We use mean absolute error (MAE) and Spearman rank correlation between the true and predicted performances.
Results. Table 3 shows the main results. Uniform random sampling, together with direct evaluation with the corresponding annotated labels, yields 3.03%p MAE and .652 rank correlation at 100 samples.
Figure 6: True and estimated accuracy on ImageNet for 50 models.
We evaluate the effectiveness of DISCO in two stages. First, we apply the model-signature approach using uniform random sampling. Then, we enhance it by selecting samples based on predictive diversity score (PDS). The results follow a similar trend. With uniform random sampling, model signatures combined with Random Forest achieve 0.86%p MAE and a rank correlation of .944, significantly outperforming the naive baseline. Incorporating PDS further improves performance, reaching 0.63%p MAE and a rank correlation of .969.
To illustrate how well the estimated performances align with the true values, we present a scatter plot in Figure 6. The high Pearson correlation coefficient of .969 indicates a strong agreement between the two.
Figure 7 shows performance across varying levels of test set reduction. The relative ranking of evaluation methods remains largely stable, except for the kNN predictor, which degrades as the number of anchor points increases. Notably, DISCO consistently outperforms all baselines, even under extreme compression with as few as 10 samples.
Figure 7: ImageNet performance estimation vs. compression rates. Mean absolute error (MAE), measured in %p difference in accuracy, and the Spearman rank correlation between the true model ranking and the estimated model ranking are shown. At 100 samples, the results are identical to Table 3. Main observations: Same as for language experiments DISCO hits a better efficiency-precision trade-off across all range of compression rates.